# Unsettling the Hegemony of Intention: Agonistic Image Generation

### v2

@@ 6. Why do different interfaces induce reflection?
To explain the observed differences in reflection under different interfaces as discussed in §5, we look at the relationship between reflection and other variables: perceived interface-level properties like appropriateness and control (§6.1), participants’ values when adding images (§6.2) and why they reject images (§6.3), and qualitative themes (§6.4).
@@ 6.1. Perceived Interface-Level Properties
Reformulative and Agonistic improve upon Baseline and Diverse in appropriateness and control ($p\leq 0.05$ for all). Indeed, in general, there is a weak but significant correlation between rethinking and both appropriateness ($r=0.22$, $p=0.02$) and control ($r=0.27$, $p=0.003$) in general. Agonistic is ranked above Diverse by $81\%$ of participants for appropriateness and $86\%$ for control; Reformulative by $85\%$ for appropriateness and $92\%$ for control. Although Diverse cannot be distinguished significantly from Baseline by appropriateness, it can by control — Diverse has lower control than Baseline ($p\ll 0.01$), meaning that rewriting user prompts to encourage diversity results in a noticeable feeling of decreased control among participants. That being said, Agonistic and Reformulative are not themselves distinguishable by control or appropriateness ratings ($p\gg 0.05$). Additionally, participants find Agonistic’s possible interpretations more interesting than Reformulative’s suggested reformulations ($p\ll 0.05$).
6.2. Image-Level Values: Why users add images
Figure 5. Distribution of values cited for images added with each interface (discussed in §6.2; Table 6).
Figure 6. Breakdown of intents by value cited when adding image (discussed in §6.2); Table 8.
@@ 6.2. Image-Level Values: Why users add images
We also explore whether different values may serve as a relevant explanatory variable for types of reflection (intents). Values code for why participants add images, or in other words the reasons that participants invoke when selecting images. The values in our coding ontology are as follows: 1) Realism: the image represents how the user thinks the world actually is; 2) Familiarity: the image fits the user’s assumptions, regardless of whether the user thinks they are representative of how the world actually is; 3) Diversity: the image portrays an underrepresented aspect of the prompt that the user believes is normatively important to include; 4) Aesthetics: the user otherwise likes how the image looks. The distribution of values across interfaces and intents is shown in Fig. 6 and 6.
@@
Agonistic is the only interface for which Diversity is invoked more frequently than Familiarity. In all interfaces except Agonistic, Familiarity is invoked most frequently, followed by Realism, Diversity, and Aesthetics. In contrast, for Agonistic, participants invoke Diversity most frequently, followed by Realism, Familiarity, and Aesthetics. Interestingly, compared to Baseline, participants invoke Diversity less frequently for Diverse ($-2\%$), but more frequently for Reformulative ($+12\%$) and Agonistic ($+19\%$).
@@
Diversity is associated with the highest rate of reflection. The majority of images across values are added with Direct intent. Among values, Aesthetics  is associated with the highest highest proportion of Direct intent ($96\%$), followed by Familiarity($95\%$), Realism ($89\%$), and Diversity ($75\%$). Diversity is associated with the highest degrees of reflection, with $7\%$ of images added with Reminder, $16\%$ of images added with Expansion, and $2\%$ of images added with Challenge. Since Agonistic also has the highest proportion of Diversity values, these findings suggest that Diversity serves as a strong explanatory variable for reflection.
6.3. Image-Level Values: Why users reject images
@@
We carry out a qualitative analysis of image rejection in lieu of a quantitative analysis because the unit of rejection is not as well-defined as adding images.
@@
Participants reject images due to factual inaccuracies, especially when interacting with Diverse. P7 said that racially diverse pictures of ‘The Chinese Communist Revolution’ “looked like modern people were slapped into the wrong time period,” and P28 experienced “pushback” against racially diverse pictures of “The Signing of the Declaration of the Independence.” In other interfaces, participants similarly cited lack of factuality as a reason for rejecting images. For instance, P2 reacted “Middle class people don’t dine out very often, and especially not at fancy places,” when Baseline generated an image of people at an upscale restaurant for the prompt “A middle class person.”
@@
Participants also reject images because of an excessive adherence to factuality. P26, generating images of “A Syrian refugee” noted that Agonistic “presented more of what I know to be true, […] the harsh reality of it. […] The other interfaces […] were more instrumental in helping me expand beyond that.P7 shared that they preferred Reformulative over Agonistic because Reformulative did a better job “highlighting all different aspects of [their prompt], instead of just the realism images.” Such comments suggest that users do value an some level of diversity in generated images that may extend beyond what they consider to be strictly factual.
@@
Participants reject images due to lack of familiarity, even where they acknowledge that an image may be factual. When using Agonistic, participants may believe interpretations to be factual (being extracted from Wikipedia) but still reject them because they are unfamiliar. For example, with the prompt “An Arab person”, P25 reacted to an image of a Sudanese Arab person from Agonistic by remarking, “Oh, Sudanese Arab […] I think that’s an identity that I haven’t represented here that I would like to.” They later decided to reject the image by providing the following reasoning: “It just does not resonate with me. I guess I’m also not familiar with Sudanese Arabs as much.
@@ 6.4. Qualitative Themes
We present several themes from qualitative analysis further explaining the differences in reflection observed in §5.
Text helps users contextualize, interpret, and decide on images. Throughout the user studies, we find that participants prefer interfaces that help them contextualize and interpret images (Reformulative and Agonistic over Baseline and Diverse), even though ultimately they are only adding the images (and not the accompanying text) to the collage. Firstly, contextualizing content is cognitively stimulating: “[the] suggestion feature… by nature makes your brain think about alternatives” (P24); it “I wouldn’t have thought of [these perspectives] without the text” (P1). Participants feel that it is meaningful and important to interpret the image, which is mediated by thinking about the relevant text. P1 said that the generated images “adhere to the prompt but you have to figure out the image yourself”. P15 said of Reformulative: “I like with this interface. It gives me a lot more of an explanation. It kind of tells a story that I might not know just by looking at it… I think that it could help me kind of change what I’m looking for.
‘Authenticity’ situates diversity in an engaged political context. Although both Diverse and Agonistic end up displaying demographically diverse images to users, we find overwhelmingly that participants prefer the type of diversity offered by Agonistic to that offered by Diverse due to a greater perception of authenticity. Diverseis not only perceived as factually inaccurate (§6.3), but nontransparent, alienating, and ungenuine. P17 said they couldn’t “get a clear idea from anything” from the interface. P5 reacted when racially and gender diverse World War II soldiers generated by Diverse: “I find them not just […] historically inaccurate, I think it gives narratives that some people will not dig deeper into, and […] can lead to a lot of misconceptions down the line of colorblind racism.” In contrast to Diverse, participants report that the perceived authenticity of Agonistic engages them in the political context of image generation (§2.2). P10 liked that Agonistic was “more grounded in reality” — “you’re like, ‘yeah, real people exist.’P28 described Agonistic: “[it] changed my perception of what the prompt itself was […] but [Diverse] feels like it’s […] trying to change our perception of the event itself that I was already thinking of […Agonistic] tries to bring in diverse perspectives and increase representation of other groups […] in a way that felt more authentic and can change people’s assumptions in a healthier way.
@@ 7.1. The inefficacy of inauthentic diversity
Though it is common to conceive of diversity and factuality as conflicting values, our findings suggest that context is more important for authentic diversity. Some participants preferred Diverse to Baseline because they felt that the demographic diversity of Diverse was appropriate and either more true to representing the subject or more familiar to their experience. For example, in response to images of a diverse crowd worshiping Jesus, P6 said,“It’s nice having […] diverse people, [be]cause I think that’s just how it is.” Nonetheless, as discussed in §6.3, many participants echoed concerns about conflicts between diversity and factuality — diversity’s “factuality tax” (Wang et al., 2023) — that also surfaced in the online reaction to the Gemini case (Milmo and Kern, 2024; Robertson, 2024). Yet notions of factuality are seldom so straightforward (§2.1). For example, when creating a collage for “the women’s suffrage movement”, P18 was hesitant to include images of non-White women because many American suffragette leaders espoused racist views, but changed their mind after engaging with an interpretation from Agonistic highlighting the role that Black women played in leading their own suffragette organizations. Here, adding the image did not obscure the facts of historical racism but highlighted other relevant aspects in the complex history of the suffragette movement. Thus although the “factuality tax” is perceived as a necessary cost of diversity, it is not clear that image generation must always be thought of along a strict dichotomy between factuality and diversity. Given that images are situated in a social context, developers should turn attention from a narrow focus on generating only the right kinds of images to also generating the right kinds of conte(x/n)t — something that users in our study found valuable (§6.4) — which can give rise to new ways of seeing and understanding the prompt.
@@
Several participants bring up aspects of a less discussed angle with Diverse’s approach to diversity: its authenticity. Taking a definition that “to say that something is authentic is to say that it is what it professes to be, or what it is reputed to be, in origin or authorship” (Varga and Guignon, 2023), Diverse can be critiqued not only as unfactual but also insufficiently aligned with the underlying goals of diversity—such as respecting and representing all groups of people. In §6.4, we documented how situating diversity in a political context encourages user perception of authenticity, as further exemplified by P18’s reflection about the suffragette movement mentioned above. These findings support the idea that the “factuality tax” is an overly restrictive framework to conceptualize authentic diversity, which should be understood neither in terms of diversity for the sake of diversity nor diversity rigidly constrained by factuality. Moreover, our findings that participants invoke Diversity at higher rates when adding images with Agonistic  ($45\%$ of images) than Diverse ($24\%$ of images) suggest that agonistic pluralism offers a superior framework for actualizing authentic diversity by engaging the user with relevant political context.
@@ 7.2. Engaging with users experientially
Our study suggests that something is missing in a picture in which user engagement with interfaces is driven by the interface’s factuality (realism, consistency) and diversity: the role of familiarity. As discussed in §6.2 and §6.3, many images are accepted and rejected because they are un/familiar rather than because they are un/factual or un/diverse. If we want to steer users towards desired ideals like critical reflection, interfaces should treat them not solely as judges for factuality and diversity, but as people who engage with images on the basis of familiarity and experience. The design of Reformulativeleans into giving users aestheticized and familiar experiences, possibly at the expense of reflection: “ if there was a change that I wanted to implement in these pictures, but I just didn’t have the words for it, I think [Reformulative]… gave the options. I didn’t have to… think about what I was looking for. I could just see examples of what I was looking for” (P19). Meanwhile, Agonistic focused not only on producing factual and diverse outputs via Wikipedia-grounded interpretation sources, but also in explicitly engaging with and challenging users’ experience and sense of familiarity. Our findings therefore further illuminate how core principles of agonistic pluralism, which emphasizes the “crucial role played by passions and affects in securing allegiance to democratic values.” (Mouffe, 2000).
@@ 7.3. Limitations and Future Work
Interface-level. Wikipedia entries are overtly sites of discourse and negotiation, and prone to political, cultural, and other forms of bias (Hube, 2017; Umarova and Mustafaraj, 2019). Even though our discovery process attempts to find varying views on subjects, this bias may filter into the process in a way that hinders agonistic user engagement. Future work might better understand the way that knowledge communities (e.g. Wikipedia, Britannica (Greenstein and Zhu, 2014)) connect with user interfaces that seek to expose users to the social and discursive processes (e.g., Wikipedia page edit history) by which information is produced. Future work on agonistic principles in generative AI interfaces might consider more explicit and iterative representations of discourse (e.g., graph-based) rather than Agonistic’s summarizing and linear presentation of discourse/interpretations.
Study-level. Although we tried to mitigate these factors, our study may be prone to novelty and participant response bias (Dell et al., 2012). Our study favored lower scale (29 participants) abut more high-resolution (1-hour interviews, collecting self-reported measures, task artifacts, qualitative responses, etc.) examinations of user experiences. Additionally, our sample population was recruited from university students in the United States, all from a single university. Future work might systematically examine how larger-scale populations engage with agonism in generative (image) AI interfaces.
8. Coda: Bricks and Windows
We introduced agonistic pluralism as a challenge to the ‘hegemony of intention’ in image generation interface design, demonstrating that an interface built on agonistic principles engaged individuals with competing interpretations in the salient discourse, making them more critically reflective. Gilles Deleuze and Felix Guattari delightfully noted that “A concept is a brick. It can be used to build a courthouse of reason. Or it can be thrown through the window” (Deleuze and Guattari, 1987). Instead of obsessing over “structurally sound” interfaces that perfectly embody capital-I Intention, capital-F Factuality, and/or capital-D Diversity, maybe our interfaces should give users some bricks to throw at the windows — they were always brittle anyways, bordering on stale, ready to be broken. Perhaps the view outside will be clearer.

### v3

@@ 6. Why do interfaces induce reflection?
To explain the observed differences in reflection under different interfaces as §5, we examine the relationship between reflection and other variables: perceived interface-level properties like appropriateness and control (§6.1), values for adding images (§6.2), values for rejecting images (§6.3), and qualitative themes (§6.4).
@@ 6.1. Perceived Interface-Level Properties
Reformulative and Agonistic surpass Baseline and Diverse in appropriateness and control ($p\leq 0.05$ for all) (Fig. 3). There is generally a weak but significant correlation between rethinking and both appropriateness ($r=0.22$, $p=0.02$) and control ($r=0.27$, $p=0.003$). Agonistic is ranked above Diverse by $81\%$ of participants for appropriateness and $86\%$ for control, while Reformulative is ranked above Diverse by $85\%$ of participants for appropriateness and $92\%$ for control among participants. Although the difference between Diverse and Baseline is not significant for appropriateness, it is for control: participants report lower control with Diverse than with Baseline ($p\ll 0.01$), suggesting that rewriting user prompts to increase diversity results in a noticeable feeling of decreased control. The difference between Agonistic and Reformulative is not significant for control or appropriateness ratings ($p\gg 0.05$), but participants find Agonistic’s interpretations more interesting than Reformulative’s suggestions ($p\ll 0.05$).
6.2. Why Users Add Images
@@
We also explore whether “value” codes may serve as a relevant explanatory variable for types of reflection (“intent” codes). As described in §4.3, “values” code for why participants add images, or more specifically the reasons that participants invoke when adding images. The values in our coding ontology are: 1) Realism: the image represents how the user thinks the world actually is; 2) Familiarity: the image fits the user’s assumptions, regardless of whether the user thinks they are representative of how the world actually is; 3) Diversity: the image portrays an underrepresented aspect of the prompt that the user believes is normatively important to include; 4) Aesthetics: the user likes how the image looks.
@@
Agonistic is the only interface for which Diversity is invoked more frequently than Familiarity (Fig. 6). In all interfaces except Agonistic, Familiarity is invoked most frequently, followed by Realism, Diversity, and Aesthetics. For Agonistic, participants invoke Diversity most frequently, followed by Realism, Familiarity, and Aesthetics. Compared to Baseline, participants invoke Diversity less frequently for Diverse ($-2\%$), but more frequently for Reformulative ($+12\%$) and Agonistic ($+19\%$), suggesting that Diverse is not an effective way to portray human diversity.
@@
Diversity is associated with the highest rate of reflection (Fig. 7). For all values, the majority of images are added with Direct intent. Aesthetics is associated with the highest proportion of Direct intent ($96\%$), followed by Familiarity ($95\%$), Realism ($89\%$), and Diversity ($75\%$). Diversity is associated with the highest degrees of reflection, with $7\%$ of images added with Reminder, $16\%$ of images added with Expansion, and $2\%$ of images added with Challenge of intent. Since Agonistic also has the highest proportion of Diversity values, these findings suggest that Diversity serves as a strong explanatory variable for reflection.
Figure 6. Distribution of value codes per interface (§6.2; Table 7).
Figure 7. Breakdown of co-occurrence between intent codes (x-axis) and value codes (y-axis) (§6.2; Table 8).
6.3. Why Users Reject Images
@@
We carry out a qualitative analysis of image rejections in lieu of a quantitative analysis because the unit of rejection is not as well-defined as added images.
@@
Participants reject images due to factual inaccuracies, especially when interacting with Diverse. Using Diverse, P7 said that racially diverse pictures of “The Chinese Communist Revolution” “looked like modern people were slapped into the wrong time period,” and P28 experienced “pushback” against racially diverse pictures of “The Signing of the Declaration of the Independence,” similar to public pushback in the Gemini controversy. Participants similarly cited lack of factuality as a reason for rejecting images in other interfaces, though to a lesser extent.
@@
Participants also reject images because of an excessive adherence to factuality. With the prompt “A Syrian refugee,” P26 noted, “[Agonistic] presented more of what I know to be true, […] the harsh reality of it. […] The other interfaces […] were more instrumental in helping me expand beyond that.” P7 preferred Reformulative over Agonistic because Reformulative did a better job ‘‘highlighting all different aspects of [the prompt], instead of just the realism images.” These comments suggest that participants value some level of diversity in generated images beyond what they consider to be strictly factual.
@@
Participants reject images due to lack of familiarity, even where they acknowledge that an image may be factual. When using Agonistic, some participants believed interpretations to be factual (being based on Wikipedia) but still rejected them because they are unfamiliar. For example, with the prompt “An Arab person,” P25 reacted to an image of a Sudanese Arab person from Agonistic: “Sudanese Arab […] that’s an identity that I haven’t represented here that I would like to.” However, they later decided to reject the image: “It just does not resonate with me. I guess I’m also not familiar with Sudanese Arabs as much.
@@ 6.4. Qualitative Themes
We present several themes from qualitative analysis that further characterize differences in reflection observed in §5.
Text helps users contextualize, interpret, and decide on images. We find that participants prefer interfaces that help them contextualize and interpret images (Reformulative and Agonistic over Baseline and Diverse), even though ultimately they are only adding the images (and not the accompanying text) to the collage. Contextualizing content is cognitively stimulating, with P24 noting that the “suggestion feature […] makes your brain think about alternatives” and P1 saying that they “wouldn’t have thought of [these images] without the text.” Participants report that text is important for mediating user interpretations of images. P1 said that the generated images “adhere to the prompt but you have to figure out the image yourself.” P15 said of Reformulative, “I like with this interface, it gives me a lot more of an explanation. It […] tells a story that I might not know just by looking at it [and] could help me kind of change what I’m looking for.
“Authentic” diversity is situated in broader political contexts. Although both Diverse and Agonistic display demographically diverse images to users, participants overwhelmingly prefer the diversity of Agonistic to that of Diverse due to a greater perception of authenticity. Participants perceive Diverse as not only factually inaccurate (§6.3), but also opaque, alienating, and disingenuous. P17 said they couldn’t “get a clear idea from anything” from Diverse, while P5 reacted to racially and gender diverse World War II soldiers from Diverse in the following way: “I find them not just […] historically inaccurate, I think it gives narratives that some people will not dig deeper into, and […] can lead to a lot of misconceptions down the line of colorblind racism.
In contrast, participants report that the perceived authenticity of Agonistic engages them in the political context of image generation. P10 liked that Agonistic was “more grounded in reality […] you’re like, ‘yeah, real people exist.’P28 shared, “[Agonistic] changed my perception of what the prompt itself was […] but [Diverse] feels like it’s […] trying to change our perception of the event itself that I was already thinking of […Agonistic] tries to bring in diverse perspectives and increase representation of other groups […] in a way that felt more authentic and can change people’s assumptions in a healthier way.
Agonistic treads the line between appropriate and inappropriate. By design, Agonistic often presents users with interpretations that are initially perceived as discomforting, inappropriate, or even offensive but may become appropriate after reflection. However, Agonistic’s interpretations may sometimes remain inappropriate even after reflection. P4 said, “[Agonistic] was giving me the most appropriate, but it was also giving me the most inappropriate [outputs] at the same time […It] would sometimes hit a home run, sometimes wildly off track.P16 said that Agonistic was “good good and bad bad […] I like the specificity [of Agonistic]. I just think the specificity was kind of whacked out.” These comments reflect the dual nature of agonistic design which by nature will tread a thin line between appropriate and inappropriate content.
@@ 7.1. The Inefficacy of Inauthentic Diversity
Since Diverse was loosely modeled after the reported behavior of the Gemini case, it provides an interesting point of analysis on the question of diversity in image generation interfaces. On one hand, some participants prefer Diverse to Baseline because they feel that the demographic diversity of Diverse is either more appropriate for representing the subject or more familiar to their experience. In response to images of a diverse crowd worshiping Jesus, P6 said,“It’s nice having […] diverse people, ’cause I think that’s just how it is.” On the other hand, as discussed in §6.3, many participants echo concerns about conflicts between diversity and factuality—diversity’s “factuality tax” (Wang et al., 2023)—that also surfaced in the online reaction to the Gemini case (Milmo and Kern, 2024; Robertson, 2024).
@@
Nonetheless, participant reactions to Agonistic also challenge conventional understandings of the tradeoffs between factuality and diversity. For example, when creating a collage for “the women’s suffrage movement,” P18 was hesitant to include images of non-White women because many American suffragette leaders espoused racist views, but changed their mind after engaging with an interpretation from Agonistic that highlighted the role of Black women in leading their own suffragette organizations. Here, adding the image did not obscure the facts of historical racism but highlighted other relevant aspects in the complex history of the suffragette movement. Thus although the “factuality tax” is perceived as a necessary cost of diversity, it is not clear that the two are always opposed to one another. Given that images are situated in sociopolitical context, developers might turn attention from a narrow focus on generating only the right kinds of images to also generating the right kinds of context—something that study participants found valuable (§6.4)—which can give rise to new ways of seeing and understanding the prompt.
@@
These and other reactions point to authenticity as a crucial yet under-discussed part of the complex interplay between diversity and factuality. Taking a definition that “[t]o say that something is authentic is to say that it is what it professes to be […] in origin or authorship” (Varga and Guignon, 2023), Diverse seems not only unfactual but also insufficiently aligned with the underlying broader goals of diversity: that is, representing and respecting underrepresented populations in ways that matter to them. In §6.4, we documented how situating diversity in a political context encourages user perception of authenticity, as further exemplified by P18’s reflection on the suffragette movement above. Moreover, our findings that participants invoke Diversity at higher rates when adding images with Agonistic than Diverse6.2) suggest that agonistic pluralism offers a superior framework for actualizing authentic diversity by engaging users with this political context rather than imposing top-down interpretations upon them. These findings further support the idea that the “factuality tax” is an overly restrictive framework for conceptualizing authentic diversity. Authentic diversity should be understood as neither diversity for the sake of diversity nor diversity rigidly constrained by factuality, but as the difference of interpretations that emerges out of real human political conflict.
@@ 7.2. Engaging with Users Experientially
Our findings highlight another component of user engagement missing from the factuality-diversity dichotomy: the role of familiarity. As discussed in §6.2 and §6.3, participants often accept and reject images because they are un/familiar rather than because they are un/factual or un/diverse. To steer users towards desired ideals like critical reflection, interfaces should treat them not solely as judges for factuality and diversity, but as people who engage with images on the basis of familiarity and experience. The design of Reformulative, for example, leans into giving users aestheticized and familiar experiences, possibly at the expense of reflection: “if there was a change that I wanted to implement in these pictures, but I just didn’t have the words for it, [Reformulative] gave the options. I didn’t have to […] think about what I was looking for. I could just see examples of what I was looking for” (P19). Meanwhile, Agonistic focuses not only on producing factual and diverse outputs via Wikipedia-grounded interpretations, but also explicitly engages and challenges users’ experiential sense of familiarity. This reflects yet another core principle of agonistic pluralism, which emphasizes the “crucial role played by passions and affects in securing allegiance to democratic values” (Mouffe, 2000).
7.3. The Sponteneity of Reflection
Despite the differences in reflection highlighted above, participants generally experienced at least some level of reflection across all interfaces (as demonstrated by Table 1), indicating that reflection is a highly subjective, personal experience. Indeed, while Baseline was uninspiring for many—“I typed what I thought and I got what I thought I would get” (P21)—it was a highly reflective experience for others—“I have a lot of other potential images in my mind of it now, and I don’t know what is […] actually true” (P28). Although reflection often occurs in planned ways (i.e., as a result of design elements we introduced with the intention of inducing reflection), it also often emerges in spontaneous yet personally significant ways as encounters with the Other (Ye et al., 2024; Halbert and Nathan, 2015; Levinas, 1969; Lacan, 2006).
@@ 7.4. Limitations and Future Work
Wikipedia entries are overtly sites of discourse and negotiation, and prone to political, cultural, and other forms of bias (Hube, 2017; Umarova and Mustafaraj, 2019; Greenstein and Zhu, 2018). Even though the Agonistic generation workflow is designed to output a diverse range of interpretations, this bias may nonetheless filter into the process in a way that hinders agonistic user engagement. Future work might seek to explore ways to design AI interfaces that allow users to engage more deeply with the discursive processes like Wikipedia edit history by which these online communities produce knowledge. Similarly, future work on agonistic design might also consider more explicit and iterative representations of discourse (e.g., graph-based), rather than Agonistic’s summarizing and linear presentation of discourse/interpretations.
In addition, we recognize that agonistic design may not be suited to all applications. We have argued that agonistic pluralism is particularly useful at challenging dominant social assumptions, but for this reason Halbert and Nathan (2015) also caution that critical reflection is less helpful for marginalized populations who already regularly face discomforting experiences in their everyday lives. Likewise, since our sample population was recruited from students at a single university, our study may be prone to our study may be prone to novelty and participant response biases (Dell et al., 2012). For example, university students may have already been predisposed to critical reflection from classroom experiences, and students at an urban institution may have already been more receptive to diverse forms of representation. Since our study favored lower-scale but highly-detailed examinations of user experiences, our results may therefore be limited in generalizability to other populations. We hope that this paper sparks further explorations of agonistic design in other contexts to better understand its limits and effectiveness at encouraging critical reflection for other populations.
8. Coda: Recentering Conflict in Diversity
From image generation interfaces emerges a political world in which individuals negotiate over the meanings of images. We introduced agonistic pluralism as a challenge to the “hegemony of intention” in current image generation interface design, demonstrating that an interface built on agonistic principles can engage individuals in this political world, making them more reflective. In her seminal essay “Situated Knowledges,” Donna Haraway uses the metaphor of vision to call for the need to “attach the objective to our theoretical and political scanners in order to name where we are and are not.” “The moral is simple,” Haraway writes: “only partial perspective promises objective vision(Haraway, 1988). By engaging users with the larger discourse they enter into upon prompting an image generation model, our interface aims to activate users’ “theoretical and political scanners” to see beyond what they intended. Agonistic interfaces surface a multitude of partial perspectives for a more “objective” user experience of image generation—more engaged with the ideas, tensions, and imaginations of real people around them.
Ultimately then, our study highlights the need for recentering conflict in diversity—not just comfortable diversity that “doesn’t make a difference,” but unsettling diversity that challenges our partial perspectives. Why is it important to care about the demographic diversity of doctors, CEOs, and Founding Fathers? Not because these cases fall under some objective standard of diversity, but because they are the products of “real acts of force” that have shaped ongoing human conflicts over representation, both in the semiotic and political sense (Melamed, 2014). Contrary to conventional narratives of diversity as a harmonious ideal, agonistic pluralism recognizes that diversity in its most authentic form stems from genuine and irreconcilable conflicts over power and representation. If such disagreements are indeed a fact of life as Mouffe suggests, then the radically democratic response for technologists is not to rush to placate user intentions by avoiding sociopolitical controversies, but rather to confront confront controversy head-on by repoening spaces where democratic conflict can take place.
Acknowledgements.
This project was partially funded by Sony Faculty Innovation Award.